OpenStudy (anonymous):

How do you solve this using the power rule? f(z) = (z^2 + 1)/3z

OpenStudy (anonymous):

what do you mean? are you wanting to take a derivative?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so rewrite...\[\frac{ \left( z^2+1 \right) }{ 3z }=\frac{ \left( z^{-1} \right)\left( z^2+1 \right) }{ 3 }=\frac{ z+z^{-1} }{ 3 }\]

OpenStudy (anonymous):

or if you like...\[\frac{ 1}{ 3 }\left( z+z^{-1} \right) \]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

you're welcome